OpenStudy (anonymous):

Find the equation of a circle that has its center at (-4, 3) and has a radius of 5. (x + 4) ^2 + (y - 3) ^2 = 25 (x - 4) ^2 + (y - 3) ^2 = 25 (x + 4) ^2 + (y + 3) ^2 = 25 (x - 4) ^2 + (y + 3) ^2 = 25

jimthompson5910 (jim_thompson5910):

hint: (x-h)^2+(y-k)^2=r^2 is a circle with center (h,k) and radius r

OpenStudy (anonymous):

so the last one

OpenStudy (jdoe0001):

\(\bf (x-{\color{green}{ h}})^2+(y-{\color{purple}{ k}})={\color{olive}{ r}}^2\qquad center\ ({\color{green}{ h}}, {\color{purple}{ k}})\quad radius = {\color{olive}{ r}}\)

jimthompson5910 (jim_thompson5910):

(x - 4) ^2 + (y + 3) ^2 = 25 is really (x - 4) ^2 + (y - (-3)) ^2 = 5^2 so h = 4 k = -3 r = 5 Center: (4,-3) Radius: 5

jimthompson5910 (jim_thompson5910):

So the center is off for choice D

OpenStudy (jdoe0001):

\(\bf (x-{\color{green}{ h}})^2+(y-{\color{purple}{ k}})={\color{olive}{ r}}^2\qquad center\ ({\color{green}{ -4}}, {\color{purple}{ 3}})\quad radius = {\color{olive}{ 5}}\)